OpenStudy (anonymous):

Using the following equation, find the center and radius: x2 − 4x + y2 + 8y = −4

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

okay, you need to complete the square on both x and y to get it in form (x-h)^2 + (y-k)^2 = r^2 and then r will be the radius, center at (h,k)

OpenStudy (anonymous):

ok so what will I haft to do next

OpenStudy (whpalmer4):

we have x^2-4x as terms in x, so we take half of the x term's coefficient, square it, and add that to both sides, giving us [x^2 - 4x + (-4/2)^2] + y^2 + 8y = -4 + (-4/2)^2 [x^2 - 4x + 4] + y^2 + 8y = -4 + 4 now we can rewrite the x portion as (x-2)^2 because (x-2)^2 = x^2 - 4x + 4 (x-2)^2 + y^2 + 8y = 0 now you repeat the process on the y portion

OpenStudy (anonymous):

how would I repeat it? please help I am very confused

OpenStudy (whpalmer4):

(x-2)^2 + [y^2 + 8y] = 0 we want to rewrite the portion in the brackets as (y-k)^2 for some value of k to do that, we take half of the value of the coefficient of the y term (not the y^2 term), square it, and add it to both sides of the equation (thus preserving the equation's correctness, but putting it in a more convenient form for us remember (y-k)(y-k) = y^2 -ky -ky + k^2 = y^2 -2ky + k^2 we have 8y and we need it to equal -2ky 8y = -2ky divide both sides by y 8 = -2k divide both sides by -2 -4 = k so we can write (y-k)^2 = (y-4)^2 = y^2 - 4y - 4y + 16 = y^2 - 8y + 16 the only problem is we don't have a 16 floating around in our left hand side of the equation, so we have to create one by adding 16 to both sides. (x-2)^2 + [y^2 + 8y + 16 ] = 0 + 16 now we can rewrite the bracketed portion as (y+4)^2 because (y+4)^2 = y^2 +8y +16 (x-2)^2 + (y+4)^2 = 16

OpenStudy (whpalmer4):

now let's compare that with our standard formula for the circle: (x-h)^2 + (y-k)^2 = r^2 the two are equal if h = 2, k = -4, r = sqrt(16) = 4 agreed?

OpenStudy (whpalmer4):

so that means the center is at (h,k) = (2, -4) the radius is r = 4

OpenStudy (anonymous):

so will the answer be thisThe center is located at (2, −4), and the radius is 4